OpenStudy (anonymous):

PLEASE HELP<3 :0 <31. What is the value of n so that the expression x^2+12x+n is a perfect square trinomial? 2. What are the solutions of the equation x^2+4x=96? 3. What are the solutions of the equation x^2+14x=-130? 4. A rocket launched from atop a 105-foot cliff with an initial velocity of 156 ft\s. The height of the rocket above the ground at a time t is given by h=-16t^2 +156t+105. When will the rocket hit the ground after it is launched? round to the nearest tenth of a second.

OpenStudy (anonymous):

Can u help @rose21

OpenStudy (rose21):

i am sorry i dont know this i am not good in math sorry :P

OpenStudy (anonymous):

Its ok:) u kn ow anyone who can?

OpenStudy (rose21):

@kmacdaddy @stonewoods

OpenStudy (anonymous):

@fallonbanks

OpenStudy (anonymous):

@osx

OpenStudy (anonymous):

@yellowlegoguy99

OpenStudy (anonymous):

Plz Anyone:(

OpenStudy (yellowlegoguy99):

For the last one... h=–16t2+156t+105 hits the ground when h = 0 solve 0=–16t2+156t+105

OpenStudy (yellowlegoguy99):

3rd one... x^2 + 14x = –130 x^2+14x+130=0 Solution=(-7+-9i)^2

OpenStudy (anonymous):

1st term of the trinomial is a perfect square 3rd term of the trinomial is a perfect square middle term is twice the product of the square root of the 1st and 3rd terms so the 3rd term (n) is 36

OpenStudy (anonymous):

THANKS GUYS:)

OpenStudy (anonymous):

Thanks @osx and number 4 is either , 4.9 s, 9.8 s, 0.6 s, 10.4 s

OpenStudy (anonymous):

i only have #1 that right... i still need 2, 3,4

OpenStudy (anonymous):

10.4s

OpenStudy (anonymous):

for #4

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

-7 + 9i and -7 -9i number 3

OpenStudy (anonymous):

number 2 -12 and 8

OpenStudy (anonymous):

for #3 i have the option of 2,-6 10, -13, 2,16 or no solution

OpenStudy (anonymous):

and for 2 i have the choice of 2, -48 -12,8 -2,48 12-8

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

your awesome

OpenStudy (anonymous):

and 2?